OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (shamim):

the equation of a line is \[y=mx+c\]

OpenStudy (shamim):

m= slope of line

OpenStudy (shamim):

so slope of ur line is \[m=2\]

OpenStudy (shamim):

if the slope of a perpendicular line is\[m _{1}\]

OpenStudy (anonymous):

y = 2x + 6 y = −x + 3 y = −x + 6 y = 2x + 3

OpenStudy (shamim):

then \[mm _{1}=-1\]

OpenStudy (shamim):

\[2m _{1}=-1\]

OpenStudy (shamim):

\[m _{1}=-\frac{ 1 }{ 2 }\]

OpenStudy (shamim):

so the equation of ur perpendicular line will b\[y-3=-\frac{ 1 }{ 2 }(x-6)\]